# Functional Brain-to-Brain Transformation with No Shared Data

### v1

@@ Abstract
Combining Functional MRI (fMRI) data across different subjects and datasets is crucial for many neuroscience tasks. Relying solely on shared anatomy for brain-to-brain mapping is inadequate. Existing functional transformation methods thus depend on shared stimuli across subjects and fMRI datasets, which are often unavailable. In this paper, we propose an approach for computing functional brain-to-brain transformations without any shared data, a feat not previously achieved in functional transformations. This presents exciting research prospects for merging and enriching diverse datasets, even when they involve distinct stimuli that were collected using different fMRI machines of varying resolutions (e.g., 3-Tesla and 7-Tesla). Our approach combines brain-to-brain transformation with image-to-fMRI encoders, thus enabling to learn functional transformations on stimuli to which subjects were never exposed. Furthermore, we demonstrate the applicability of our method for improving image-to-fMRI encoding of subjects scanned on older low-resolution 3T fMRI datasets, by using a new high-resolution 7T fMRI dataset (scanned on different subjects and different stimuli).
@@ 1 Introduction
Functional MRI (fMRI) has emerged as a powerful non-invasive method for measuring human brain activity in real-time, allowing researchers to observe which areas of the brain are involved in different functions and behaviors [1, 2, 3, 4, 5]. However, the time-consuming and expensive data acquisition process of fMRI often leads to limited data availability. Moreover, inter-subject variability in brain structure and functionality poses additional challenges in analyzing fMRI data from multiple subjects, as different individuals may exhibit distinct brain activity patterns even in response to similar stimuli. To address these limitations, functional brain-to-brain transformation has been proposed. Functional brain-to-brain transformation involves learning inter-subject transformations that map the fMRI signals of one subject to the fMRI signals of another. Those transformations allow for the use of data collected from multiple subjects, thereby enriching the available data and enabling comparisons of brain activity across individuals and groups.
@@
While the visual cortex exhibits similarities in topography and organization of functional regions across humans, there are fine-grained functional differences between individuals [6, 7, 8, 9]. This variability poses challenges in comparing and mapping fMRI activation between individuals. As such, traditional anatomical alignment methods [10, 11, 12, 13] have limited functional predictive power across subjects [10, 14, 15, 16]. To address this issue, functional alignment methods have been developed to match the functional behaviors of different individuals by learning relationships between brain activity patterns [14, 15, 17, 18, 19]. There are two main approaches in functional alignment: the first maps subjects into a high-dimensional shared space, with the most well-known method being Hyper-alignment [14, 20, 21, 22], and the second transforms directly between fMRI of two subjects using a conversion model [15]. However, training these methods requires a large amount of “shared data” namely, fMRI obtained by showing the same stimuli to different subjects. This limits the number of available training examples, and makes it impossible to learn transformations across subjects which have no shared data. The requirement for shared data thus undermines the ability to combine information from multiple different fMRI datasets, accumulated over the years from a wide variety of visual stimuli from different individuals.
@@
In this paper, we introduce for the first time, an approach for computing functional brain-to-brain transformation without any shared data. Figure 1 formulates the problem setting. Our approach (illustrated in Figure 2), employs “visual encoders”, which are trained individually on subject-specific data, to encode images into subject-specific fMRI. Our method effectively uses visual encoders to predict fMRI of images not seen by a subject. This allows to use also “non-shared data” (which refers to images seen by only one subject), and “external images” (which refers to images not seen by any of the subjects, i.e. arbitrary images without fMRI recordings). To the best of our knowledge, we are the first to employ visual encoders to compute brain-to-brain transformations, thereby exploiting both “non-shared” and “external” data. By leveraging this approach, we significantly enrich the transformation training data. Notably, we demonstrate the capability to compute functional brain-to-brain transformations without any shared images, even when the fMRI data was collected using different fMRI scanners (e.g., 7T vs 3T). Furthermore, our results show that we can improve one subject’s image-to-fMRI encoder by utilizing another subject’s encoder, even in the absence of shared data. In particular, we demonstrate the ability to enhance the image-to-fMRI encoding of a subject trained on a low-resolution 3-Tesla fMRI dataset (e.g., “GOD” dataset [23]) by incorporating another subject’s encoder trained on a new 7-Tesla fMRI dataset (e.g., NSD [24]) using our proposed method.
@@
  • Enhancing the encoding of one subject using data from another high-quality subject (both within and across datasets, e.g., utilizing a subject from a new 7-Tesla dataset to enhance a subject from a low-resolution 3-Tesla dataset).
@@ 2 Overview of the Approach
Our approach aims to learn brain-to-brain transformations between subjects (within and across datasets) when subjects were exposed to different stimuli, possibly acquired by different fMRI machines (see Figure 1). We first explain how using visual encoders allows generating “corresponding” fMRI data between different subjects without any shared stimuli. This “corresponding” fMRI data is used to train the transformation. We further demonstrate the effectiveness and applicability of our approach for improving the encoding of one subject using high-quality data from another subject (scanned on a different stimuli dataset and on a different fMRI scanner).
Brain-to-brain transformation with no shared data:
@@
Our approach involves the use of “visual encoders”, which encode images into fMRI signals. These are modeled via deep neural networks, trained on subject-specific image-fMRI pairs. Visuals encoders were shown to obtain impressive accuracy in forecasting a subject’s fMRI responses to novel images [25, 26, 27, 28, 29, 30]. We first individually train a visual encoder for each subject, utilizing their own respective image-fMRI data (see Figure 2.a). Then, we use the encoder training approach of [30]. Note that such encoder training does not require for the two subjects to have experienced the same stimuli; they can belong to entirely distinct image-fMRI datasets.
@@
The functional brain-to-brain transformation $T$ is trained to map the fMRI patterns of Subject1 to their corresponding counterparts in Subject2. The straightforward approach to train the transformation network is to transform an fMRI scan from one subject to match the corresponding fMRI scan of the second subject, and then evaluate its error w.r.t the recorded fMRI (e.g., using the mean square error loss). However, this approach relies on the presence of shared data, and is moot when such shared data is unavailable. Leveraging the predictive capability of the pre-trained encoders for predicting fMRI responses of images, we can generate corresponding fMRI patterns for “non-shared” and “external” images. Given an fMRI scan $\mathcal{F}_{1}$ of Subject1, linked to an image $\mathcal{I}_{1}$ that only this subject saw (as shown in Figure 2.b on the left side), we can approximate the corresponding fMRI pattern for Subject2 by inputting image $\mathcal{I}_{1}$ into the encoder ${E}_{2}$ designed for Subject2. We then optimize the learned transformation T using the fMRI loss between the transformed fMRI $T(\mathcal{F}_{1})$ and the encoded fMRI ${E_{2}(\mathcal{I}_{1})}$, aiming to minimize differences between the two predicted fMRIs ($||T(\mathcal{F}_{1})-E_{2}(\mathcal{I}_{1})||$). A similar process can be carried out using the fMRI pattern from an image observed solely by Subject2, comparing it against the encoding of the same image using the encoder for Subject1 ($||T({E}_{1}(\mathcal{I_{2}})-\mathcal{F}_{2})||$). Furthermore, we can incorporate “external” images (depicted in Figure 2.b on the right side). These images are natural images that were never seen by any of the subjects, and thus do not have any fMRI recordings. We feed these “external” images into the encoders of both subjects, subsequently transforming the encoded fMRI pattern of Subject1 and comparing it with the encoded fMRI pattern of Subject2 ($||T({E}_{1}(\mathcal{I}))-E_{2}(\mathcal{I})||$). This provides “infinitely” many training data for self-supervised training of the brain-to-brain transformation $T$.
@@
Figure 2: An overview of the approach - Estimating functional brain-to-brain transformations with no shared data: (a) Initially, we train separate visual encoders for each subject using its own image-fMRI pairs dataset. Each subject’s data can potentially come from entirely different image datasets and different fMRI scanners. (b) The transformation $T$ is then computed by training a single linear layer to correctly transform fMRI patterns of Subject1 to those of Subject2. We employ the subject-specific pre-trained encoders to predict fMRI responses both for “non-shared” images (i.e, images seen by only one subject), and for “external” images (any natural image never seen by any of the subjects). The large set of predicted fMRIs is then used to train the transformation network $T$.
Improving Image-to-fMRI encoding of one subject using high-quality fMRI of another subject:
@@
To showcase the power of our method, we employ it to enhance the image-to-fMRI encoder of one subject using another subject with superior data quality or more extensive examples. For simplicity, we term this technique the “teacher-student” method, where the subject with higher quality data plays the role of the “teacher”, and the poorer-quality subject is the “student”. This “teacher-student” approach is applicable to subjects of different quality within the same dataset, and more interestingly, it can be employed across different datasets and machines using the above-described method.
@@ 3 Results
Figure 3: The fMRI datasets used in our experiments: (i) “Vim-1” dataset [31] which features grey-scale images and their corresponding 3-Tesla fMRI recordings. (ii) the “Generic Object Decoding” (GOD) dataset [32] which comprises natural ImageNet images with their corresponding 3-Tesla fMRI recordings. (iii) the “Natural Scenes Dataset” (NSD), a new 7-Tesla dataset [24] with 8 subjects, each of whom viewed 1000 shared images and 9000 unique subject-specific images.
@@
We have experimented with three prominent fMRI datasets (see Figure 3): (i) the “vim-1” dataset, which features around 2000 grey-scale images and their corresponding 3-Tesla fMRI recordings for 2 subjects [31, 33, 34]; (ii) the “Generic Object Decoding” (GOD) dataset, which comprises of 1300 pairs of natural images from ImageNet with 3-Tesla fMRI recordings for 5 subjects [32, 35, 36]; and (iii) the “Natural Scenes Dataset” (NSD), a new 7-Tesla dataset with 8 subjects, each having around 1000 shared images and 9000 unique subject-specific images [24] (all in all - 73000 different images).
3.1 Results of Brain-to-Brain Transformation with No Shared Data
@@
This section presents results which substantiate two central claims: (i) Functional transformations can be computed without relying on shared data, yielding significantly better results than anatomical mapping; (ii) Harnessing “non-shared” and “external” images (either alongside or instead of shared data), leads to significantly improved transformations compared to those estimated with shared data.
To validate these claims, we use the NSD dataset. This dataset offers a relatively limited number of shared data (1000 shared images) and a substantial number of non-shared data (9000 images per subject), rendering it well-suited for illustrating our claims. Given that we possess both non-shared and shared data, we can exclusively train a transformation model using non-shared data and then assess its performance on a dedicated evaluation set of shared data. We evaluate the quality of the learned transformation $T$ in 2 ways: (i) Quantitatively, we measure the Pearson correlation between the predicted fMRI (i.e., the fMRI obtained by transforming the source subject’s fMRI to the target subject’s space) and the actual fMRI of the target subject; (ii) Qualitatively, we assess the content of the transformed fMRI through an fMRI decoder (mapping fMRI to images), and visually inspect the quality of the decoded images. This decoding is achieved by combining 2 fMRI-to-image decoding methods [30] and [37] (for further details, please refer to the Supplementary Material).
@@
Figure 4: Evaluating brain-to-brain transformation with no shared data: (a) Quantitative evaluation: The graph presents the mean Pearson’s r correlation averaged over all 56 transformations of all possible pairs of subjects in the NSD dataset, with SEM (standard error of the mean) error bars. We compare the resulting transformations when trained only with non-shared data (solid red line) to three other transformations: an anatomical mapping to a common brain space (cyan dotted line), a transformation trained with only 200 shared examples (dotted blue line), and a transformation trained with 700 shared examples (dotted grey line). The remaining 300 “shared data” serve as “test data” for assessing and comparing the quality of all the learned transformations. The x-axis of the graph represents the number of non-shared examples available for training, where our transformation model is the only one that can use non-shared examples. (b) Qualitative evaluation: This figure provides a visual comparison of image decoding when using our transformation model versus anatomical alignment method (the only 2 possible methods for handling non-shared data). As can be seen in the diagram, the transformation is used to map the fMRI activity of Subject1 to predict the fMRI activity of Subject2. Then we utilize the pre-trained decoder [30] of Subject2 (trained on his mutually exclusive data), together with a novel diffusion model decoding approach [37] to generate images from the transformed fMRI. The figure displays the original image seen by Subject1, along with two reconstructed images post-fMRI transformation from Subject1 to Subject2: Once using anatomical mapping between the subjects, and once using our transformation model $T$ trained without any shared data.
Quantitative evaluation: The graph in Figure 4.a presents the mean Pearson’s r correlation averaged over all 56 transformations of all possible pairs of subjects in the NSD dataset. We compare the resulting transformations when trained only with non-shared data (solid red line) to three other transformations: an anatomical mapping to a common brain space (cyan dotted line), a transformation trained with only 200 shared examples (dotted blue line), and a transformation trained with 700 shared examples (dotted grey line). The remaining 300 “shared data” serve as “test data” for assessing and comparing the quality of all the learned transformations. The x-axis of the graph represents the number of non-shared examples available for training, where our transformation model is the only one that can use non-shared examples. When there are no shared examples, we can only compare our method to anatomical alignment. As seen from the graph, our approach performs significantly better than anatomical alignment, even with small amounts of non-shared examples ($df=56$, $t=18.57$, $p<1e-20$, $d=2.48$; up to 75% improvement when training with as little as 200 non-shared examples). Furthermore, our method, when utilizing only non-shared examples, yields comparable or even better results than transformation models trained with shared data. As we increase the number of non-shared examples, the transformation model’s performance continues to improve, surpassing models trained with only 200 and 700 shared examples. Although shared data is more beneficial when there is a low number of examples, 700 non-shared examples (M=0.407, SD=0.07) perform better than 200 shared examples (M=0.385, SD=0.07, $df=56$, $t=12.08$, $p<5e-17$, $d=1.61$), and 1600 non-shared examples (M=0.4349, SD=0.07) perform better than all 700 shared examples (M=0.429, SD=0.07, $df=56$, $t=3.32$, $p=2e-3$, $d=0.44$). This is because, with sufficient examples, the encoder used for our training of the transformation is robust enough to compensate for the absence of shared examples. We provide specific transformations with mean Pearson’s r correlation values in Figure LABEL:S1 in the Supplementary Material. Figures LABEL:S2 and LABEL:S3 further demonstrate that combining non-shared examples with available shared examples leads to better results than using shared examples alone.
@@
Qualitative evaluation: Figure 4.b further provides a visual comparison of visual decoding when using our transformation model versus anatomical alignment method (the only two possible methods for handling non-shared data). The transformation is used to map the fMRI activity of Subject1 to predict the fMRI activity of Subject2. We then utilize a pre-trained decoder [30] of Subject2 (trained on its unique data), together with a novel diffusion model decoding approach [37] to generate images from the transformed fMRI. The figure displays the original image seen by Subject1, along with two reconstructed images post fMRI transformation from Subject1 to Subject2: Once using anatomical mapping between the subjects, and once using our transformation model $T$ trained without any shared data. This demonstrates that our transformation produces fMRI predictions with more relevant and semantically meaningful information for reconstructing images. This visual comparison underscores that our transformation, trained without shared data, generates significantly superior outcomes compared to anatomical mapping, the only technique used to-date when shared data is absent. For additional reconstructions and transformations, please refer to Figures LABEL:S4 and LABEL:S5 in the Supplementary Material.
3.2 Improving Image-to-fMRI Encoding of one Subject using another Higher-Quality Subject
@@
We next demonstrate how our transformation method allows one subject from a high quality fMRI dataset (the “teacher”) to improve the image-to-fMRI encoding capabilities of another subject from a low quality fMRI dataset (the “student”). Normally there are no shared images across different fMRI datasets, in which case our transformation method is needed. This applies also when both subjects are from the same dataset, but one subject has higher fMRI quality than the other.
Figure 5: Improving encoding of a subject from GOD 3-Tesla dataset (“student”) using another high-quality subject (“teacher”): This figure compares the quality of the “student” baseline encoder (which is trained solely on the student’s own fMRI data) to the student encoder obtained when incorporating also data from the “teacher” during training. In each plot, the grey line represents the student baseline encoder model, the purple line symbolizes the encoder improvement obtained when using a “teacher” from the same dataset (e.g., Subject4 is the highest-quality subject in the GOD dataset). The orange line corresponds to the encoder improvement achieved by utilizing a superior subject from another dataset (NSD), which has more scanned examples and higher fMRI resolution (7-Tesla machine). The quality of the resulting encoders is assessed through the mean Pearson correlation of all predicted fMRI voxels. The “number of training examples” refers to the number of the student’s own image-fMRI pairs of examples used for training the “student” encoder, whereas the “teacher” subject (drawn from either GOD or NSD dataset) has been trained using all its own available examples and remains unchanged during the training of the student encoder. (a) Results for an individual student subject (Subject1). (b) Aggregated results of all five subjects in GOD dataset (displaying mean values and SEM error bars). (c) The distribution of Pearson correlations across the fMRI voxels, comparing the baseline student encoder against the improved student encoder when trained with a teacher from the NSD dataset. This is evaluated using varying numbers of training examples (300, 600, and 1200) from the student’s own fMRI data. The x-axis signifies the Pearson correlation values, while the y-axis represents the density of fMRI voxels corresponding to each correlation value.

### v2

@@ Abstract
Combining Functional MRI (fMRI) data across different subjects and datasets is crucial for many neuroscience tasks. Relying solely on shared anatomy for brain-to-brain mapping is inadequate. Existing functional transformation methods thus depend on shared stimuli across subjects and fMRI datasets, which are often unavailable. In this paper, we propose an approach for computing functional brain-to-brain transformations without any shared data, a feat not previously achieved in functional transformations. This presents exciting research prospects for merging and enriching diverse datasets, even when they involve distinct stimuli that were collected using different fMRI machines of varying resolutions (e.g., 3-Tesla and 7-Tesla). Our approach combines brain-to-brain transformation with image-to-fMRI encoders, thus enabling to learn functional transformations on visual stimuli to which subjects were never exposed. Furthermore, we demonstrate the applicability of our method for improving image-to-fMRI encoding of subjects scanned on older low-resolution 3T fMRI datasets, by using a new high-resolution 7T fMRI dataset (scanned on different subjects and different stimuli).
@@ 1 Introduction
Functional MRI (fMRI) has emerged as a powerful non-invasive method for measuring human brain activity , allowing researchers to observe which areas of the brain are involved in different functions and behaviors [1, 2, 3, 4, 5]. However, the time-consuming and expensive data acquisition process of fMRI often leads to limited data availability. Moreover, inter-subject variability in brain structure and functionality poses additional challenges in analyzing fMRI data from multiple subjects, as different individuals may exhibit distinct brain activity patterns even in response to similar stimuli. To address these limitations, functional brain-to-brain transformation has been proposed. Functional brain-to-brain transformation involves learning inter-subject transformations that map the fMRI signals of one subject to the fMRI signals of another. Those transformations allow for the use of data collected from multiple subjects, thereby enriching the available data and enabling comparisons of brain activity across individuals and groups.
@@
While the visual cortex exhibits similarities in topography and organization of functional regions across humans, there are fine-grained functional differences between individuals [6, 7, 8, 9]. This variability poses challenges in comparing and mapping fMRI activation between individuals. As such, traditional anatomical alignment methods [10, 11, 12, 13] have limited mapping prediction accuracy across subjects [10, 14, 15, 16]. To address this issue, functional alignment methods have been developed to match the functional behaviors of different individuals by learning relationships between brain activity patterns [14, 15, 17, 18, 19, 20, 21]. There are two main approaches in functional alignment: the first maps subjects into a high-dimensional shared space, with the most well-known method being Hyper-alignment [14, 22, 23, 24], and the second transforms directly between fMRI of two subjects using a conversion model [15]. However, training these methods requires a sufficient amount of “shared data” namely, fMRI obtained by showing the same stimuli to different subjects. This limits the number of available training examples, and makes it impossible to learn transformations across subjects which have no shared data. The requirement for shared data thus undermines the ability to combine information from multiple different fMRI datasets, accumulated over the years from a wide variety of visual stimuli from different individuals.
@@
In this paper, we introduce for the first time, an approach for computing functional brain-to-brain transformation without any shared data, and show its effectiveness for fMRI responses to visual stimuli. Figure 1 formulates the problem setting. Our approach (illustrated in Figure 2), employs “visual encoders”, which are trained individually on subject-specific data, to encode images into subject-specific fMRI. Our method effectively uses visual encoders to predict fMRI of images not seen by a subject. This allows to use also “non-shared data” (which refers to images seen by only one subject), and “external data” (which refers to images not seen by any of the subjects, i.e. arbitrary images without fMRI recordings). To the best of our knowledge, we are the first to employ visual encoders to compute brain-to-brain transformations, thereby exploiting both “non-shared” and “external” data. By leveraging this approach, we significantly enrich the transformation training data (i.e., the number of examples used for learning the functional mapping between subjects). Notably, we demonstrate the capability to compute functional brain-to-brain transformations without any shared images, even when the fMRI data was collected using different fMRI scanners (e.g., 7T vs 3T). Furthermore, our results show that we can improve one subject’s image-to-fMRI encoder by utilizing another subject’s encoder, even in the absence of shared data. In particular, we demonstrate the ability to enhance the image-to-fMRI encoding of a subject trained on a low-resolution 3-Tesla fMRI dataset (e.g., “GOD” dataset [25]) by incorporating another subject’s encoder trained on a new 7-Tesla fMRI dataset (e.g., NSD [26]) using our proposed method.
@@
  • Enhancing the encoding of one subject using higher quality data from another subject (both within and across datasets, e.g., utilizing a subject from a new 7-Tesla dataset to enhance a subject from a low-resolution 3-Tesla dataset).
@@ 2 Overview of the Approach
Our approach aims to learn Brain-to-Brain (B2B) transformations between subjects (within and across datasets), when the subjects were exposed to different stimuli, possibly acquired by different fMRI machines (see Figure 1). We first explain how using visual encoders allows generating “corresponding” fMRI data between different subjects without any shared stimuli. By “corresponding” fMRI we mean synthesizing the fMRI responses of two different subjects, as if they were exposed to the same image stimulus. This “corresponding” fMRI data is used to train the B2B-transformation across different subjects. We further demonstrate the effectiveness and applicability of our B2B-transformations for improving the Image-to-fMRI encoding of one subject using higher-quality fMRI data from another subject (scanned on a different stimuli dataset and on a different fMRI scanner).
2.1 Brain-to-Brain (B2B) transformation with no shared data
@@
Our approach involves the use of “visual encoders”, which encode images into fMRI signals. These are modeled via deep neural networks, trained on subject-specific image-fMRI pairs. Image-to-fMRI encoders were shown to obtain impressive accuracy in forecasting a subject’s fMRI responses to novel images [27, 28, 29, 30, 31, 32]. We first individually train a visual encoder for each subject, utilizing their own respective image-fMRI data (see Figure 2.a). Note that such encoder training does not require for the two subjects to have experienced the same stimuli; they can belong to entirely distinct image-fMRI datasets.
@@
The functional brain-to-brain transformation $T$ is trained to map the fMRI patterns of Subject1 to their corresponding counterparts in Subject2. The straightforward approach to train the transformation network is to transform an fMRI scan from one subject to match the corresponding fMRI scan of the second subject, and then evaluate its error w.r.t the recorded fMRI (e.g., using the mean square error loss). However, this approach relies on the presence of shared data, and is moot when such shared data is unavailable. Leveraging the predictive capability of the pre-trained encoders for predicting fMRI responses of images, we can generate corresponding fMRI patterns for “non-shared” and “external” images. For example, given a “non-shared” image $\mathcal{I}_{Ns_{1}}$ seen only by Subject1, and Subject1’s fMRI scan on that image, $\mathcal{F}_{1}(\mathcal{I}_{Ns_{1}})$ (as shown in Figure 2.b on the left side), we can approximate the corresponding fMRI pattern for Subject2 by inputting image $\mathcal{I}_{Ns_{1}}$ into the encoder ${E}_{2}$ designed for Subject2. We then optimize the learned transformation T (modeled by a single linear layer) using the fMRI loss between the transformed fMRI $T(\mathcal{F}_{1}(\mathcal{I}_{Ns_{1}}))$ and the encoded fMRI ${E_{2}(\mathcal{I}_{Ns_{1}})}$, aiming to minimize differences between the two predicted fMRIs: ($||T(\mathcal{F}_{1}(\mathcal{I}_{Ns_{1}}))-E_{2}(\mathcal{I}_{Ns_{1}})||$). A similar process can be carried out using the fMRI scan of a “non-shared” image observed solely by Subject2, $\mathcal{F}_{2}(\mathcal{I}_{Ns_{2}})$, comparing it against the transformed fMRI encoding of the same image using the encoder of Subject1: ($||T({E}_{1}(\mathcal{I}_{Ns_{2}}))-\mathcal{F}_{2}(\mathcal{I}_{Ns_{2}}))||$). Furthermore, we can incorporate “external” images (depicted in Figure 2.b on the right side). These images are natural images that were never seen by any of the subjects, and thus do not have any fMRI recordings. We feed these “external” images into the encoders of both subjects, subsequently transforming the encoded fMRI pattern of Subject1 and comparing it with the encoded fMRI pattern of Subject2: ($||T({E}_{1}(\mathcal{I}_{Ext}))-E_{2}(\mathcal{I}_{Ext})||$). This provides “infinitely” many training data for training of the brain-to-brain transformation $T$. While real fMRI scans are generally more reliable than encoded fMRIs predicted from images, the primary advantage of using external images lies in their infinite supply, which is particularly beneficial when the number of available fMRI-image pairs is small.
@@
Figure 2: An overview of the approach - Estimating functional brain-to-brain transformations with no shared data: (a) We first independently train a visual encoder for each subject using its own dataset of image-fMRI pairs. Each subject’s data can potentially come from entirely different image datasets and different fMRI scanners. (b) We employ the subject-specific pre-trained encoders to predict fMRI responses both for “non-shared” images (i.e, images seen by only one subject), and for “external” images (any natural image never seen by any of the subjects). The transformation $T$ is then computed by training a single linear layer to correctly transform fMRI patterns of Subject1 to those of Subject2 using the large set of predicted fMRIs. The lock symbol signifies that the encoder weights are fixed and not updated while training the transformation (in this scenario).
2.2 Improving Image-to-fMRI Encoding via B2B Transformation
@@
To showcase the power of our method, we employ it to enhance the image-to-fMRI encoder of one subject using another subject with superior data quality or more examples. For simplicity, we term this technique the “teacher-student” method, where the subject with higher quality data plays the role of the “teacher”, and the poorer-quality subject is the “student”. This “teacher-student” approach is applicable to subjects of different quality within the same dataset, and more interestingly, it can be employed across different datasets and machines using the above-described method.
@@ 3 Results
Figure 3: The fMRI datasets used in our experiments: (i) “Vim-1” dataset [33] which features grey-scale images and their corresponding 4-Tesla fMRI recordings. (ii) the “Generic Object Decoding” (GOD) dataset [34] which comprises natural ImageNet images with their corresponding 3-Tesla fMRI recordings. (iii) the “Natural Scenes Dataset” (NSD), a new 7-Tesla dataset [26] with 8 subjects, each of whom viewed 1000 shared images and 9000 unique subject-specific images. This resulted in a total of 73,000 different natural images (taken from the COCO dataset).
@@
We have experimented with three prominent fMRI datasets (see Figure 3): (i) the “vim-1” dataset [33, 35, 36], which contains around 2000 grey-scale images (as opposed to color images) and their corresponding 4-Tesla fMRI recordings for 2 subjects; (ii) the “Generic Object Decoding” (GOD) dataset [34, 37, 38], which comprises of 1300 pairs of natural images from ImageNet with 3-Tesla fMRI recordings for 5 subjects; and (iii) the “Natural Scenes Dataset” (NSD) [26], a new 7-Tesla dataset with 8 subjects, each of whom viewed 9000 unique subject-specific images, in addition to $\sim$1000 shared images viewed by all subjects. This resulted in a total of 73,000 different natural images (taken from the COCO dataset). See 5.1 for details about train/val/test splits for each dataset.
Results of Brain-to-Brain (B2B) Transformation with No Shared Data
@@
This section presents results which substantiate two central claims: (i) Functional B2B transformations can be computed without relying on shared data, yielding significantly better results than anatomical mapping; (ii) Harnessing a large number of “non-shared” and/or “external” images (either alongside or instead of shared data), leads to improved B2B transformations compared to those estimated when using only the small number of available shared data.
To validate these claims, we used the NSD dataset. We automatically selected 10,000 voxels that are most sensitive to visual tasks, based on their signal-to-noise ratio (SNR), which is a standard approach in the field (see Methods Section 5.2 for more details). This dataset offers a relatively limited number of shared data (1000 shared images) and a substantial number of non-shared data (9000 subject-specific images per subject), rendering it well-suited for illustrating our claims. We train a transformation model exclusively on the non-shared data, and then assess its performance on a dedicated evaluation set of shared data. For each subject we independently trained an Image-to-fMRI encoder on its own subject-specific image-fMRI pairs (see Table S2 for each encoder’s prediction correlation). These encoders are then used for training the B2B transformations. We evaluate the learned B2B transformation $T$ both quantitatively and qualitatively.
@@
Figure 4: Quantitative evaluation of brain-to-brain transformation with no shared data: (a)  The graph presents the mean Pearson’s r correlation averaged over all 56 transformations of all possible pairs of subjects in the NSD dataset, with SEM (standard error of the mean) error bars. We compare the resulting transformations when trained only with non-shared data (solid red line) to three other transformations: an anatomical mapping to a common brain space (cyan dotted line), a transformation trained with only 200 shared examples (dotted blue line), and a transformation trained with 700 shared examples (dotted grey line). The remaining 300 “shared data” serve as “test data” for assessing and comparing the quality of all the learned transformations. The x-axis of the graph represents the number of non-shared examples available for training, where our transformation model is the only one that can use non-shared examples. (b) The figure presents a correlation heatmap for all NSD subject pairs, using solely non-shared and external data.
@@
Figure 5: Qualitative evaluation of brain-to-brain transformation with no shared data: This figure provides a visual comparison of image decoding when using our transformation model versus anatomical alignment method (the only 2 possible methods for handling non-shared data). As can be seen in the diagram, the transformation is used to map the fMRI activity of Subject1 to predict the fMRI activity of Subject2. Then we utilize the pre-trained decoder [32] of Subject2 (trained on his mutually exclusive data), together with a novel diffusion model decoding approach [39] to generate images from the transformed fMRI. The figure displays the original image seen by Subject1, along with two reconstructed images post-fMRI transformation from Subject1 to Subject2: Once using anatomical mapping between the subjects, and once using our transformation model $T$ trained without any shared data.
Quantitative evaluation: We measure the Pearson correlation between the predicted fMRI (i.e., the fMRI obtained by transforming the source subject’s fMRI to the target subject’s space), and the actual fMRI of the target subject. The graph in Figure 4.a presents the mean Pearson’s r correlation averaged over all 56 transformations of all possible pairs of subjects in the NSD dataset. We compare the resulting transformations when trained with non-shared data together with 6000 external images (solid red line) to three other transformations: an anatomical mapping to a common brain space (cyan dotted line), a transformation trained with only 200 shared examples (dotted blue line), and a transformation trained with 700 shared examples (dotted grey line). The remaining 300 “shared data” serve as “test data” for assessing and comparing the quality of all the learned B2B transformations. The x-axis of the graph represents the number of non-shared examples available for training, where our transformation model is the only one that can use non-shared examples. When there are no shared examples whatsoever, we can only compare our method to anatomical alignment. As seen from the graph, our approach performs significantly better than anatomical alignment, even with a small number of non-shared examples (paired samples permutation test with $N=56$, $p<0.00001$, $d=2.48$) with up to 75% improvement when training with as little as 200 non-shared examples). Furthermore, our method, when utilizing only non-shared examples, yields comparable or even better results than transformation models trained with shared data. As we increase the number of non-shared examples, the transformation model’s performance continues to improve, surpassing models trained with only 200 and 700 shared examples. Although shared data is more beneficial when there is a low number of examples, 700 non-shared examples (M=0.407, SD=0.07) perform better than 200 shared examples (M=0.385, SD=0.07, $N=56$, $p<0.00001$, $d=1.61$), and 1600 non-shared examples (M=0.4349, SD=0.07) perform better than all 700 shared examples (M=0.429, SD=0.07, $N=56$, $p<0.001$, $d=0.44$). This is likely because, with sufficient examples, the encoder used for our training of the transformation is robust enough to compensate for the absence of shared examples.
In Figure 4.b we present a correlation heatmap of all NSD subject pairs trained using only non-shared and external data. Additional results in the Supplementary Material show transformations for specific subjects (Figures S4 and S3), and demonstrating that combining non-shared examples with shared examples yields better performance than using shared examples alone (Figures S5 and S6). Furthermore, the capability to learn B2B transformations without shared data enables us to train transformations across different datasets with differing fMRI resolutions. In the Supplementary Material (Figure S2), we provide more detailed quantitative results of these B2B transformations (for any pair of subjects), to showcase this ability. This capability is later utilized to enhance the encoding of one dataset using another.
Qualitative evaluation: We further assess the quality of the B2B-transformed fMRI through an fMRI-to-Image decoder, and visually inspect the quality of the decoded images. This decoding is achieved by combining 2 fMRI-to-Image decoding methods – [32] and [39] (for more details, please refer to Section 5.6). Figure 5.b provides a qualitative visual comparison of images decoded from the transformed fMRI – once after using our B2B transformation model, and once after anatomically mapping the fMRI (the only two possible methods for handling non-shared data). The transformation is used to map the fMRI activity of NSD Subject1 to predict the fMRI activity of NSD Subject2. We then utilize the decoder of Subject2 (pre-trained on its unique subject-specific data), to generate images from the transformed fMRI. The figure displays the original image seen by Subject1, along with two reconstructed images post fMRI transformation from Subject1 to Subject2: One obtained by using anatomical mapping between the two subjects, and one using our B2B-transformation model $T$ trained without any shared data. This demonstrates that our B2B-transformation produces fMRI predictions with more relevant and semantically meaningful information for reconstructing images. This visual comparison underscores that our transformation, trained without shared data, generates significantly superior transformed fMRIs compared to anatomical mapping (the only available option to-date). For additional reconstructions and transformations, please refer to Figures S7 and S8 in the Supplementary Material.
Improving Image-to-fMRI Encoding via B2B Transformation
@@
Ours B2B transformation allows one subject from a high quality fMRI dataset (the “teacher”) to improve the image-to-fMRI encoding capabilities of another subject from a low quality fMRI dataset (the “student”). Normally there are no shared images across different fMRI datasets, in which case our B2B transformation method is needed. This applies also when both subjects are from the same dataset, but one subject has higher fMRI quality than the other. We assess the significance and utility of the enhanced encoder through 3 evaluations: (i) Pearson correlation between predicted and measured fMRI; (ii) the ability to perform accurate Image Retrieval via the encoded fMRI; and (iii) the influence of the encoded fMRI on Image Classification accuracy.